# Labels Matter More Than Models: Quantifying the Benefit of Supervised Time Series Anomaly Detection

### v1

@@ C Full Results of STAND Sensitivity Analysis
Figures A9, A10, and A11 illustrate the impact of the model dimension ($d_{\text{model}}$), the number of TEM layers, and the window size on STAND’s performance across the five datasets, respectively. Overall, the model’s performance improves (for almost all metrics) as the model dimension increases. Conversely, performance tends to degrade (for both point-wise and event-level metrics) as the number of TEM layers and the temporal window size increase.
Figures A9, A10, and A11 also present the impact of $d_{\text{model}}$, the number of TEM layers, and the window size on STAND’s performance across the four PSM subsets. Generally, the trend of model performance change is consistent with that observed on the five real-world datasets. Specifically, when the model dimension increases, the performance improves on three subsets (SP2–SP4), while the performance change on SP1 is relatively marginal. Conversely, when the number of TEM layers and the temporal window size increase, performance slightly decreases on the SP2–SP4 subsets, but no clear trend of performance change is observed on SP1.

### v2

@@ C Full Results of STAND Sensitivity Analysis
Figures A10, A11, and A12 illustrate the impact of the model dimension ($d_{\text{model}}$), the number of TEM layers, and the window size on STAND’s performance across the five datasets, respectively. Overall, the model’s performance improves (for almost all metrics) as the model dimension increases. Conversely, performance tends to degrade (for both point-wise and event-level metrics) as the number of TEM layers and the temporal window size increase.
Figures A10, A11, and A12 also present the impact of $d_{\text{model}}$, the number of TEM layers, and the window size on STAND’s performance across the four PSM subsets. Generally, the trend of model performance change is consistent with that observed on the five real-world datasets. Specifically, when the model dimension increases, the performance improves on three subsets (SP2–SP4), while the performance change on SP1 is relatively marginal. Conversely, when the number of TEM layers and the temporal window size increase, performance slightly decreases on the SP2–SP4 subsets, but no clear trend of performance change is observed on SP1.